### 12 Apr 2026 v1

# Two Lemmas on the Fiber-wise Holomorphicity in Complex Algebraic Geometry

Abstract
Motivated by Hartogs’ theorem on separate holomorphicity, we establish two rigidity results in complex geometry.
First, we prove a lemma on regularity for weak solutions of algebraic differential equations, namely, we show that a fiber-wise distributional $L^{2}$ solution to an algebraic linear differential equation automatically upgrades to a global holomorphic solution on the total space, assuming the equation admits a compatible $L^{2}$ distributional solution along a transverse compact complex submanifold that maps birationally to the base.
Second, for a compact Kobayashi hyperbolic manifold fibered over $\mathbb{P}^{1}$, we prove that a continuous, fiber-wise holomorphic map of degree $1$ to a projective variety is a bi-holomorphic isomorphism, provided it is injective on a very ample hypersurface.
Keywords: Complex manifold, Hyperbolicity, Hartogs’ theorem, Birational geometry, Algebraic differential equation
1 Introduction and Background
The rigid behavior of holomorphic functions of several complex variables is one of the foundational pillars of complex analysis. A striking manifestation of this rigidity is the ability to deduce global holomorphicity from merely partial and lower-dimensional analytic data. In 1906, F. Hartogs proved the following fundamental theorem on separate holomorphicity:
Theorem 1.1 (F. Hartogs, [1]).
Let $\Delta:=\{z\in\mathbb{C}:|z|<1\}$ be the open unit disc in the complex plane $\mathbb{C}$, and let $f\colon\Delta^{n}\rightarrow\mathbb{C}$ be an arbitrary function. Suppose that for every $(a_{1},\dots,a_{n})\in\Delta^{n}$ and every $i\in\{1,\dots,n\}$, the function
$g\colon\Delta\rightarrow\mathbb{C},\quad z\mapsto f(a_{1},\dots,a_{i-1},z,a_{i+1},\dots,a_{n})$
is holomorphic. Then $f$ itself is holomorphic.
In accordance with Hartogs’ theorem, we investigate the properties of fiber-wise holomorphic mappings.
We recall that in the theory of complex analysis, the Cauchy–Fantappiè integral formula is a fundamental tool for proving holomorphicity. However, transitioning from bounded domains to projective varieties presents a profound topological difference: compact complex manifolds possess no boundary. To replace the classical role of boundary integrals in forcing interior regularity, one must employ global algebraic geometry. In this paper, we present two related rigidity lemmas demonstrating how fiber-wise analytic behavior, when anchored by transverse algebraic intersections, such as rational sections and ample divisors, automatically elevates to global algebraic regularity.
2 A Lemma on Algebraic Differential Equations
In modern partial differential equation theory, it is standard to initially obtain solutions in a weak sense, such as $L^{2}$ distributions. The following result establishes a bridge, proving that if such weak fiber-wise solutions can be anchored by a compatible $L^{2}$ distributional solution on a transverse submanifold, their moduli are perfectly locked. This geometric anchoring forces the weak data to automatically patch together, bypassing the need for a priori global uniform estimates and elevating the solution to be globally holomorphic across the entire manifold.
Theorem 2.1.
Let $X,Y$ be smooth projective varieties over $\mathbb{C}$, and $f\colon X\rightarrow Y$ a surjective morphism with simply connected generic fibers. Let $M$ be a compact complex submanifold of $X$ such that $f|_{M}$ is birational. Let $\mathscr{E}\rightarrow X$ be a holomorphic vector bundle equipped with a flat connection $\nabla$, and $\sigma$ a holomorphic section of $\Omega^{1}_{X}\otimes\mathscr{E}$ satisfying $\nabla\sigma=0$. Suppose that there exist an $L^{2}$ section $\psi_{a}$ of $\mathscr{E}|_{f^{-1}(a)}$ for every regular value $a\in Y$ of $f$ and an $L^{2}$ section $\phi$ of $\mathscr{E}|_{M}$ solving the system
$\begin{cases}(i^{*}\nabla)\phi=\sigma|_{M},\\ (i_{a}^{*}\nabla)\psi_{a}=\sigma|_{f^{-1}(a)};\end{cases}$ (1)
in the sense of distributions, where $i\colon M\rightarrow X$ and $i_{a}\colon f^{-1}(a)\rightarrow X$ are the canonical inclusions. Then there exists a holomorphic solution $\Phi$ to the equation $\nabla\Phi=\sigma$ such that $\phi=\Phi|_{M}$.
Proof.
By generic smoothness, there exists a Zariski dense open subset $U$ of $Y$, such that for every $a\in U$, the fiber $f^{-1}(a)$ is an irreducible smooth projective variety. Decomposing the connection as $\nabla=\nabla^{1,0}+\bar{\partial}$, the fiber-wise equations in system (1) imply $\bar{\partial}\phi=\bar{\partial}\psi_{a}=0$, as $\sigma$ is a $(1,0)$-form. By Weyl’s lemma on elliptic regularity applied to the $\bar{\partial}$-operator, the distributional solutions to system (1) are classical, and hence holomorphic.
Since the generic fibers of $f$ are simply connected, their fundamental groups vanish, making all flat vector bundles over $f^{-1}(a)$ trivializable. Any parallel section on $f^{-1}(a)$ is therefore uniquely determined by its evaluation at a single point without monodromy obstructions. Notice that, for a generic $a\in U$, the fiber $f^{-1}(a)$ intersects $M$ transversely. Since $f|_{M}$ is birational, there exists a Zariski dense open subset $V$ of $U$ over which $f|_{M}$ restricts to an isomorphism. By the vanishing of the obstruction class to solving the fiber-wise differential equation $(i_{a}^{*}\nabla)\psi=\sigma|_{f^{-1}(a)}$ and the holomorphic dependence on initial values, there exists a holomorphic local section $\Phi$ of $\mathscr{E}$ which solves the parametric family of complex PDEs
$\begin{cases}(i_{a}^{*}\nabla)\Phi|_{f^{-1}(a)}=\sigma|_{f^{-1}(a)},\\ \Phi|_{M}=\phi.\end{cases}$ (2)
For any $a\in V$, we have that $M$ intersects the fiber $f^{-1}(a)$ at exactly one point, rendering the initial value problem (2) well-posed. The local solution $\Phi$ thus extends to a global holomorphic section over $f^{-1}(V)$.
We now claim that the equality $\nabla\Phi=\sigma$ holds $f^{-1}(V)$. Consider the residual section $R:=-\sigma+\nabla\Phi$. Utilizing the flatness of the connection $\nabla$ alongside the closedness of $\sigma$, we deduce
$\nabla R=\nabla(\nabla\Phi-\sigma)=\nabla^{2}\Phi-\nabla\sigma=0.$
By the very construction, the restriction of $R$ to the fiber directions vanishes. Thus, for any vertical vector field $\xi$ tangent to the fibers, we have $R(\xi)=0$. Let $\eta$ be any local projectable horizontal vector field. Contracting $\nabla R$ with the bi-vector $\xi\otimes\eta$, we obtain
$(\nabla R)(\xi,\eta)=\nabla_{\xi}(R(\eta))-\nabla_{\eta}(R(\xi))-R([\xi,\eta])=0$
via Cartan’s formula for the exterior covariant derivative. Since $R$ annihilates vertical vectors, $\nabla_{\eta}(R(\xi))=0$. Because $\eta$ is projectable and $\xi$ is vertical, their Lie bracket $\xi,\eta]$ is purely vertical, implying $R([\xi,\eta])=0$ also vanishes, which reduces the equation to $\nabla_{\xi}(R(\eta))=0$, meaning the tensor contraction of $R$ with any horizontal vector field is a covariant constant along the fibers. Furthermore, when restricted to the transverse submanifold $M$, the tangent space splitting $T_{x}X=T_{x}M\oplus T_{x}(f^{-1}(a))$ forcing the tensor $R$ to be determined at $x$ by its pullbacks. Note that
$i^{*}R=(i^{*}\nabla)\Phi|_{M}-i^{*}\sigma=(i^{*}\nabla)\phi-\sigma|_{M}=0$
by the first equation in system (1). The transversality then yields that $R(\eta)$ vanishes at the intersection points of the fibers of $f$ with $M$. Since $R(\eta)$ is constant along the fibers and vanishes on $M$, it vanishes identically. Since $R(\xi)=R(\eta)=0$, we conclude that $R=0$, and hence the equation $\nabla\Phi=\sigma+R=\sigma$ holds on $f^{-1}(V)$. Because $\Phi$ satisfies $\nabla\Phi=\sigma$ globally on $f^{-1}(V)$, its local boundedness near the complement $X\setminus f^{-1}(V)$ follows from the holomorphic Poincaré lemma. Specifically, since $\nabla\sigma=0$, there exists a holomorphic local section s{s} satisfying $\nabla{s}=\sigma$ in a sufficiently small neighborhood of any point $x\in X\setminus f^{-1}(V)$. Consequently, we arrive at $\nabla(\Phi-{s})=0$ in the vicinity of $x$, which implies that $\Phi-{s}$ is a constant vector. The local boundedness of $\Phi$ is thus immediate. Consequently, the section $\Phi$ extends holomorphically across the proper analytic subset $X\setminus f^{-1}(V)$ to the entirety of $X$ by Riemann’s extension theorem, preserving the identity $\nabla\Phi=\sigma$ globally. ∎
Remark 2.2.
The assumption of an $L^{2}$ fiber-wise solution makes this lemma particularly adaptable to geometric analysis settings, allowing one to rigidly extend weak distributional limits directly without establishing global uniform continuity beforehand.
3 A Lemma on Kobayashi Hyperbolic Manifolds
We now transition from functions to mappings. In this section, we demonstrate that continuous maps from a Kobayashi hyperbolic manifold can be elevated to global isomorphisms, using a very ample hypersurface to lock the complex moduli.
Theorem 3.1.
Let $\phi\colon X\rightarrow Y$ be a continuous map of mapping degree $1$ between compact complex $n$-manifolds of equal dimension $n\geq 4$, where $Y$ is projective and $X$ is Kobayashi hyperbolic. Suppose that there exist a holomorphic fibration $f\colon X\rightarrow\mathbb{P}^{1}$ and a very ample hypersurface $H$ in $X$ such that $\phi|_{f^{-1}(t)}$ is holomorphic for every $t\in\mathbb{P}^{1}$ and $\phi|_{H}$ is injective. Then $\phi$ is a bi-holomorphic isomorphism.
Proof.
We divide the proof into three main steps: Establishing that the restrictions of $\phi$ on the fibers are finite, then proving their images are mutually disjoint, and finally upgrading the continuous map $\phi$ to a biholomorphism.
Step 1: Since $H\in\operatorname{div}(X)$ is ample, the complex manifold $X$ is projective by Kodaira’s embedding theorem. Let
$\Sigma:=\mathbb{P}^{1}-\{b_{1},\dots,b_{m}\}$
be the punctured Riemann sphere consisting of the regular values of $f$. For any $t\in\Sigma$, the fiber $X^{t}:=f^{-1}(t)$ is a smooth projective subvariety of $X$ of pure dimension n1n-1. Since $\phi^{t}:=\phi|_{X^{t}}$ is holomorphic and $Y$ is projective, Remmert’s proper mapping theorem and Chow’s theorem ensure that the image $Y^{t}:=\phi(X^{t})$ is a projective subvariety of $Y$. We first establish the injectivity of the restriction $\phi^{t}$ for any fixed $t\in\Sigma$. For the sake of contradiction, we now assume that $\phi^{t}$ is not injective.
Suppose first that $\dim(Y^{t})\leq n-3$. By Chevalley’s fiber dimension theorem, the generic fiber of $\phi^{t}$ would have dimension at least 2, and hence would contain a projective surface. Since the intersection of the ample divisor $H$ with a projective surface contains a curve in $H$, the morphism $\phi^{t}$ would contract a curve in $H$ to a point, violating the injectivity of $\phi|_{H}$. Therefore, we must have at least $\dim(Y^{t})\geq n-2$.
Suppose now that $\dim(Y^{t})=n-2$. Take a generic point $y\in Y^{t}$. Then the fiber $C$ of $\phi^{t}$ over $y$ is a projective curve. Since $H\in\operatorname{div}(X)$ is a very ample divisor, $H^{t}:=X^{t}\cap H$ is very ample in $X^{t}$, making $H^{t}\cap C$ non-empty and $\phi(H^{t})=Y^{t}$. Because $\phi|_{H}$ is injective by assumption, the mapping degree of $\phi^{t}|_{H^{t}}$ onto its image is exactly 1. By choosing $y$ to be generic, we ensure that the restriction $\phi^{t}|_{H^{t}}$ is unramified at its unique preimage $x\in H^{t}$, rendering its differential an isomorphism at $x$. As a result, the tangent space $T_{x}H^{t}$ intersects trivially with the kernel of the differential of $\phi^{t}$ at $x$, which is precisely $T_{x}C$. This proves that $C$ intersects $H^{t}$ transversally, and hence their intersection product is $H^{t}\cdot C=1$. Because $H$ is a very ample divisor, intersecting a projective curve $C$ at a single point with a multiplicity of exactly 1 forces $C$ to be rational, that is, bi-holomorphic to $\mathbb{P}^{1}$. However, as $X$ is Kobayashi hyperbolic, there exists no non-constant holomorphic mapping from $\mathbb{P}^{1}$ into $X$ by Brody’s theorem on hyperbolicity. This is a contradiction. Therefore, we have in fact $\dim(Y^{t})=n-1$, so the regular mapping $\phi^{t}\colon X^{t}\rightarrow Y$ sends its domain of definition onto a projective variety $Y^{t}$ of equal dimension n1n-1. We claim that $\phi^{t}$ is actually a finite morphism. Consider the Stein factorization $\phi^{t}=\phi^{t}_{1}\circ\phi^{t}_{2}$ of $\phi^{t}$, where $\phi^{t}_{2}\colon X^{t}\rightarrow N$ is a surjective projective morphism with connected fibers onto a normal projective variety $N$, and $\phi^{t}_{1}\colon N\rightarrow Y^{t}$ is a finite morphism. Since $\phi^{t}_{1}$ is finite and surjective, we obtain
$\dim(N)=\dim(Y^{t})=n-1.$
Because $\phi^{t}_{2}$ is a surjective morphism with connected fibers between projective varieties of the same dimension, its generic fiber is a single point, implying that $\phi^{t}_{2}$ is a birational morphism. Assume to the contrary that $\phi^{t}$ admits a fiber $F$ of positive dimension. Since $\phi^{t}_{1}$ is finite, this fiber $F$ must be contracted to a point entirely by $\phi^{t}_{2}$. Therefore, the exceptional locus of the birational morphism $\phi^{t}_{2}$ is non-empty. Now, recall from birational geometry that the exceptional locus of a proper birational morphism from a smooth projective variety is uniruled, meaning it is covered by rational curves. However, as $X$ is Kobayashi hyperbolic, by Brody’s theorem, it contains no rational curves, which leads to a contradiction. This proves that $\phi^{t}$ is indeed a finite morphism.
Step 2: Now, we prove that $Y^{t_{1}}$ and $Y^{t_{2}}$ cannot intersect for distinct $t_{1},t_{2}\in\Sigma$. Assume to the contrary that there exist points $x_{1}\in X^{t_{1}}$ and $x_{2}\in X^{t_{2}}$ for distinct $t_{1},t_{2}\in\Sigma$ such that $\phi(x_{1})=\phi(x_{2})$. Since $\phi^{t}$ is a finite surjective morphism onto its image for every $t\in\Sigma$, the images $Y^{t_{1}}$ and $Y^{t_{2}}$ are projective subvarieties of $Y$ of dimension n1n-1. By Krull’s dimension theorem, the intersection $Y^{t_{1}}\cap Y^{t_{2}}$ is a projective subvariety of $Y$ with dimension bounded below by
$\dim(Y^{t_{1}}\cap Y^{t_{2}})\geq(n-1)+(n-1)-n=n-2\geq 2.$
Consider the incidence variety
$I:=\{(a_{1},a_{2})\in X^{t_{1}}\times X^{t_{2}}\mid\phi(a_{1})=\phi(a_{2})\}.$
Because $\phi^{t_{1}}$ and $\phi^{t_{2}}$ are finite morphisms, the natural projection from $I$ onto $Y^{t_{1}}\cap Y^{t_{2}}$ has finite fibers. Therefore, $I$ is a projective subvariety of the fiber product $X^{t_{1}}\times X^{t_{2}}$ with
$\dim(I)=\dim(Y^{t_{1}}\cap Y^{t_{2}})\geq 2.$
Let $p_{1}\colon I\rightarrow X^{t_{1}}$ and $p_{2}\colon I\rightarrow X^{t_{2}}$ be the restrictions of the canonical projections onto each factor. Because $\phi^{t_{1}}$ and $\phi^{t_{2}}$ are finite morphisms, the projections $p_{1}$ and $p_{2}$ are also finite. Therefore, the pullbacks $p_{1}^{*}(H)$ and $p_{2}^{*}(H)$ are both ample divisors on the projective variety $I$. Since $\dim(I)\geq 2$, their intersection product $p_{1}^{*}(H)\cdot p_{2}^{*}(H)$ is strictly positive, ensuring the intersection of their support is non-empty. Take any $(h_{1},h_{2})\in p_{1}^{*}(H)\cap p_{2}^{*}(H)$. Then, by the very construction, we have $h_{1},h_{2}\in H$ such that $\phi(h_{1})=\phi(h_{2})$, which again contradicts the injectivity of $\phi|_{H}$. We conclude that the images $Y^{t}$ are mutually disjoint for all $t\in\Sigma$. Step 3: Because $Y$ is a smooth projective variety, and the collection $\{Y^{t}:t\in\Sigma\}$ consists of homologous effective divisors that are pairwise disjoint, the intersection product $Y^{t_{1}}\cdot Y^{t_{2}}$ vanishes for all distinct $t_{1},t_{2}\in\Sigma$. Therefore, a Zariski dense open subset of the compact complex manifold $Y$ is continuously foliated by the hypersurfaces $Y^{t}$ over $\Sigma$, yielding a continuous mapping
$\upsilon\colon\Sigma\rightarrow\operatorname{Chow}(Y),\quad t\mapsto Y^{t}$
into the Chow variety of $Y$. Because the connected components of $\operatorname{Chow}(Y)$ are all projective, we may consider the Zariski closure $V$ of the image $\upsilon(\Sigma)$ within $\operatorname{Chow}(Y)$. Since $\Sigma$ is an uncountably infinite set and the members of $\upsilon(\Sigma)$ are mutually disjoint, the image $\upsilon(\Sigma)$ consists of uncountably many distinct points. As a 0-dimensional variety over $\mathbb{C}$ is necessarily a finite set of points, the Zariski closure $V$ must have dimension $\dim(V)\geq 1$. This implies that $V$ is a projective family of effective divisors on $Y$ of strictly positive dimension. Because the condition of having empty intersections is an open condition in the Zariski topology, and $\upsilon(\Sigma)$ is by construction Zariski dense in $V$, the generic members of the projective family $V$ are also mutually disjoint. Since the effective divisors in $V$ are algebraically equivalent and generically disjoint, their intersection product vanishes. On any smooth projective variety, a positive-dimensional projective family of effective divisors with vanishing mutual intersection numbers must possess an empty base locus. Consequently, the linear equivalence classes of divisors in $V$ span a free pencil, which guarantees the existence of a surjective regular morphism $g\colon Y\rightarrow\mathbb{P}^{1}$ such that $g^{-1}(t)=Y^{t}$ for every $t\in\mathbb{P}^{1}$. Because $\phi$ maps the fibers of $f$ into the fibers of $g$, we have $g\circ\phi=f$. Since the mapping degree is multiplicative, we have
$\deg(\phi^{t})=\deg(\operatorname{id}_{\mathbb{P}^{1}})\cdot\deg(\phi^{t})=\deg(\phi)=1$
for all $t\in\mathbb{P}^{1}$. Since $Y$ is a smooth projective variety and $g\colon Y\rightarrow\mathbb{P}^{1}$ is a holomorphic fibration, Bertini’s theorem ensures that the general fiber of $g$ is smooth, and therefore normal. Thus, for a generic $t\in\Sigma$, the finite morphism $\phi^{t}\colon X^{t}\rightarrow Y^{t}$ of degree 1 onto a normal variety is an isomorphism by Zariski’s main theorem. This in particular proves that $\phi$ is injective almost everywhere.
Take an arbitrary $t\in\Sigma$. Since $X$ is a compact Kobayashi hyperbolic manifold, we have that $X^{t}$ is also Kobayashi hyperbolic. Recall from birational geometry that any rational map into a projective variety containing no rational curves is automatically a regular morphism. Therefore, the group of birational mappings from $X^{t}$ to itself coincides with $\operatorname{Aut}(X^{t})$, which is finite as
$H^{0}(X^{t};\operatorname{Der}(\mathcal{O}_{X^{t}}))=0$
by the rigidity of Kobayashi hyperbolic manifolds. Consider the relative morphism scheme $\Phi\colon\operatorname{Hom}_{\mathbb{P}^{1}}(X,Y)\rightarrow\mathbb{P}^{1}$, where $X$ and $Y$ are viewed as fibrations over $\mathbb{P}^{1}$ via $f$ and $g$ respectively. For any two finite morphisms $\varphi_{1},\varphi_{2}\colon X^{t}\rightarrow Y^{t}$ of degree 1, the composition
$(\varphi_{1})^{-1}\circ\varphi_{2}\colon X^{t}\rightarrow X^{t}$
is birational. Consequently, the set of all finite morphisms from $X^{t}$ to $Y^{t}$ of degree 1 forms a principal homogeneous space under the $\operatorname{Aut}(X^{t})$-action. Since $\operatorname{Aut}(X^{t})$ is finite, the set $Z^{t}$ of all finite morphisms of degree 1 in the fiber of $\Phi\colon\operatorname{Hom}_{\mathbb{P}^{1}}(X,Y)\rightarrow\mathbb{P}^{1}$ over any $t\in\Sigma$ is finite. Moreover, since the Zariski closure of
$Z:=\{(t,\varphi):t\in\Sigma,\varphi\in Z^{t}\}$
in $\operatorname{Hom}_{\mathbb{P}^{1}}(X,Y)$ is precisely the connected component of $\operatorname{Hom}_{\mathbb{P}^{1}}(X,Y)$ containing the finite morphism $\phi^{t}$ of mapping degree 1, it is a projective curve such that $\Phi|_{Z}$ is a branched covering. Since $\phi$ is uniformly continuous, it induces a continuous section
$\phi^{*}\colon\Sigma\rightarrow Z,\quad t\mapsto\phi^{t}$
of the branched covering $\Phi|_{Z}$. Since $\Phi|_{Z}$ admits a continuous section over $\Sigma$, it is in fact not branched over $\Sigma$. Moreover, any continuous section of an unbranched analytic covering is necessarily holomorphic, so $\phi^{*}$ is holomorphic on $\Sigma$. Because $\phi$ is holomorphic along fibers and the section $\phi^{*}$ is also holomorphic, Osgood’s lemma on separate holomorphicity guarantees that $\phi$ is holomorphic on $f^{-1}(\Sigma)$. Since $\phi$ is continuous and the singular locus $X-f^{-1}(\Sigma)$ forms a proper algebraic subset, Riemann’s extension theorem guarantees that $\phi$ is holomorphic on the entirety of $X$.
Having established that $\phi\colon X\rightarrow Y$ is a proper holomorphic mapping of degree 1 between smooth projective varieties, it remains to show it is an isomorphism. But, if $\phi$ were not an isomorphism, as it is injective almost everywhere, it would be a birational map contracting an exceptional divisor to a lower-dimensional subvariety. Recall from birational geometry that the exceptional locus of a proper birational morphism between smooth projective varieties is uniruled, which once again contradicts the hyperbolicity of $X$.
Quod erat demonstrandum. ∎
Acknowledgement
The author is deeply grateful to Weiyi Zhang for very useful and inspiring discussions. This research was completed while the author was studying at the Mathematics Institute of the University of Warwick. The author therefore would like to thank the University of Warwick for its hospitality.
Funding Information
No funding was received to assist with the preparation of this manuscript, and the author did not receive support from any organization for the submitted work.
Statements and Declarations
The author certifies that the author has no affiliations with or involvement in any other organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript. This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author and the source, provide a link to the Creative Commons licence, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons licence, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons licence and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder. Data sharing is not applicable to this article as no datasets were generated or analysed during the current study. The author hereby provides consent for the publication of the manuscript detailed above.

### 16 Apr 2026 v2

@@
-Two Lemmas on the Fiber-wise Holomorphicity in Complex Algebraic Geometry
+A Note on Two Useful Lemmas in Complex Algebraic Geometry
 Abstract
-Motivated by Hartogs’ theorem on separate holomorphicity, we establish two rigidity results in complex geometry.
-First, we prove a lemma on regularity for weak solutions of algebraic differential equations, namely, we show that a fiber-wise distributional $L^{2}$ solution to an algebraic linear differential equation automatically upgrades to a global holomorphic solution on the total space, assuming the equation admits a compatible $L^{2}$ distributional solution along a transverse compact complex submanifold that maps birationally to the base.
+We establish two useful lemmas in complex geometry.
+First, we investigate the homological properties of differential operators intertwined by the Dolbeault Laplacian on compact Hermitian manifolds.
+By utilizing the mapping cone of Koszul complexes, we prove that the analytical obstructions to solving specific overdetermined PDE systems reduce entirely to the finite-dimensional Dolbeault cohomology.
 Second, for a compact Kobayashi hyperbolic manifold fibered over $\mathbb{P}^{1}$, we prove that a continuous, fiber-wise holomorphic map of degree $1$ to a projective variety is a bi-holomorphic isomorphism, provided it is injective on a very ample hypersurface.
-Keywords: Complex manifold, Hyperbolicity, Hartogs’ theorem, Birational geometry, Algebraic differential equation
+Keywords: Dolbeault cohomology, Koszul complex, Complex manifold, Hyperbolicity
 1 Introduction and Background
@@
 is holomorphic. Then $f$ itself is holomorphic.
-In accordance with Hartogs’ theorem, we investigate the properties of fiber-wise holomorphic mappings.
-We recall that in the theory of complex analysis, the Cauchy–Fantappiè integral formula is a fundamental tool for proving holomorphicity.
-However, transitioning from bounded domains to projective varieties presents a profound topological difference: compact complex manifolds possess no boundary.
-To replace the classical role of boundary integrals in forcing interior regularity, one must employ global algebraic geometry.
-In this paper, we present two related rigidity lemmas demonstrating how fiber-wise analytic behavior, when anchored by transverse algebraic intersections, such as rational sections and ample divisors, automatically elevates to global algebraic regularity.
-2 A Lemma on Algebraic Differential Equations
-In modern partial differential equation theory, it is standard to initially obtain solutions in a weak sense, such as $L^{2}$ distributions.
-The following result establishes a bridge, proving that if such weak fiber-wise solutions can be anchored by a compatible $L^{2}$ distributional solution on a transverse submanifold, their moduli are perfectly locked.
-This geometric anchoring forces the weak data to automatically patch together, bypassing the need for a priori global uniform estimates and elevating the solution to be globally holomorphic across the entire manifold.
-Theorem 2.1.
-Let $X,Y$ be smooth projective varieties over $\mathbb{C}$, and $f\colon X\rightarrow Y$ a surjective morphism with simply connected generic fibers.
-Let $M$ be a compact complex submanifold of $X$ such that $f|_{M}$ is birational.
-Let $\mathscr{E}\rightarrow X$ be a holomorphic vector bundle equipped with a flat connection $\nabla$, and $\sigma$ a holomorphic section of $\Omega^{1}_{X}\otimes\mathscr{E}$ satisfying $\nabla\sigma=0$.
-Suppose that there exist an $L^{2}$ section $\psi_{a}$ of $\mathscr{E}|_{f^{-1}(a)}$ for every regular value $a\in Y$ of $f$ and an $L^{2}$ section $\phi$ of $\mathscr{E}|_{M}$ solving the system
-$\begin{cases}(i^{*}\nabla)\phi=\sigma|_{M},\\ (i_{a}^{*}\nabla)\psi_{a}=\sigma|_{f^{-1}(a)};\end{cases}$ (1)
-in the sense of distributions, where $i\colon M\rightarrow X$ and $i_{a}\colon f^{-1}(a)\rightarrow X$ are the canonical inclusions.
-Then there exists a holomorphic solution $\Phi$ to the equation $\nabla\Phi=\sigma$ such that $\phi=\Phi|_{M}$.
+In accordance with the spirit of rigidity exemplified by Hartogs’ theorem, this paper investigates how rigid global structures in complex geometry can be enforced by partial algebraic or homological constraints.
+We divide our exploration into two distinct but philosophically related results.
+In Section 2, we turn our attention to differential operators on compact Hermitian manifolds.
+In the study of integrable systems, operators intertwined by a Laplacian exhibit highly constrained solution spaces.
+We abstract this phenomenon using homological algebra.
+By establishing an acyclicity lemma for intertwining chain complexes, we demonstrate how the Dolbeault Laplacian acts as an invertible intertwining operator on the space of exact forms.
+This allows us to prove that the infinite-dimensional analytical obstructions to solving certain coupled PDE systems collapse entirely into the finite-dimensional Dolbeault cohomology.
+In Section 3, we transition from the rigidity of differential equations to the rigidity of mappings.
+While Hartogs’ theorem relies on boundary integrals within bounded domains, transitioning to compact projective varieties requires global algebraic geometry.
+We present a rigidity lemma demonstrating how continuous, fiber-wise analytic mappings can be automatically elevated to global bi-holomorphic isomorphisms when anchored by a transverse algebraic intersection, specifically a very ample hypersurface, on a Kobayashi hyperbolic manifold.
+2 A Lemma on Holomorphic Differential Equations
+We begin our investigation by formulating a purely algebraic statement.
+By localizing the homological complexity to a specific submodule where an intertwining operator acts as an isomorphism, we equate the cohomology of the total space with that of the quotient.
+Lemma 2.1.
+Let $0\rightarrow A\rightarrow B\rightarrow C\rightarrow 0$ be a short exact sequence of $R$-modules, and let $f,g,h\in\operatorname{End}_{R}(B)$ satisfy $fg=gh$, so that the sequence $K(B)$ of $R$-modules given by
+[…]
+is a chain complex.
+Suppose that $h(A)\subseteq g(A)=A$ and $g|_{A}$ is a monomorphism.
+Then $0\rightarrow A\rightarrow A^{\oplus 2}\rightarrow A\rightarrow 0$ forms an acyclic subcomplex $K(A)$ of $K(B)$, such that the cohomology groups of $K(C):=K(B)/K(A)$ are isomorphic to that of $K(B)$.
 Proof.
-By generic smoothness, there exists a Zariski dense open subset $U$ of $Y$, such that for every $a\in U$, the fiber $f^{-1}(a)$ is an irreducible smooth projective variety.
-Decomposing the connection as $\nabla=\nabla^{1,0}+\bar{\partial}$, the fiber-wise equations in system (1) imply $\bar{\partial}\phi=\bar{\partial}\psi_{a}=0$, as $\sigma$ is a $(1,0)$-form.
-By Weyl’s lemma on elliptic regularity applied to the $\bar{\partial}$-operator, the distributional solutions to system (1) are classical, and hence holomorphic.
-Since the generic fibers of $f$ are simply connected, their fundamental groups vanish, making all flat vector bundles over $f^{-1}(a)$ trivializable.
-Any parallel section on $f^{-1}(a)$ is therefore uniquely determined by its evaluation at a single point without monodromy obstructions.
-Notice that, for a generic $a\in U$, the fiber $f^{-1}(a)$ intersects $M$ transversely.
-Since $f|_{M}$ is birational, there exists a Zariski dense open subset $V$ of $U$ over which $f|_{M}$ restricts to an isomorphism.
-By the vanishing of the obstruction class to solving the fiber-wise differential equation $(i_{a}^{*}\nabla)\psi=\sigma|_{f^{-1}(a)}$ and the holomorphic dependence on initial values, there exists a holomorphic local section $\Phi$ of $\mathscr{E}$ which solves the parametric family of complex PDEs
-$\begin{cases}(i_{a}^{*}\nabla)\Phi|_{f^{-1}(a)}=\sigma|_{f^{-1}(a)},\\ \Phi|_{M}=\phi.\end{cases}$ (2)
-For any $a\in V$, we have that $M$ intersects the fiber $f^{-1}(a)$ at exactly one point, rendering the initial value problem (2) well-posed.
-The local solution $\Phi$ thus extends to a global holomorphic section over $f^{-1}(V)$.
-We now claim that the equality $\nabla\Phi=\sigma$ holds $f^{-1}(V)$.
-Consider the residual section $R:=-\sigma+\nabla\Phi$.
-Utilizing the flatness of the connection $\nabla$ alongside the closedness of $\sigma$, we deduce
-$\nabla R=\nabla(\nabla\Phi-\sigma)=\nabla^{2}\Phi-\nabla\sigma=0.$
-By the very construction, the restriction of $R$ to the fiber directions vanishes.
-Thus, for any vertical vector field $\xi$ tangent to the fibers, we have $R(\xi)=0$.
-Let $\eta$ be any local projectable horizontal vector field.
-Contracting $\nabla R$ with the bi-vector $\xi\otimes\eta$, we obtain
-$(\nabla R)(\xi,\eta)=\nabla_{\xi}(R(\eta))-\nabla_{\eta}(R(\xi))-R([\xi,\eta])=0$
-via Cartan’s formula for the exterior covariant derivative.
-Since $R$ annihilates vertical vectors, $\nabla_{\eta}(R(\xi))=0$.
-Because $\eta$ is projectable and $\xi$ is vertical, their Lie bracket $\xi,\eta]$ is purely vertical, implying $R([\xi,\eta])=0$ also vanishes, which reduces the equation to $\nabla_{\xi}(R(\eta))=0$, meaning the tensor contraction of $R$ with any horizontal vector field is a covariant constant along the fibers.
-Furthermore, when restricted to the transverse submanifold $M$, the tangent space splitting $T_{x}X=T_{x}M\oplus T_{x}(f^{-1}(a))$ forcing the tensor $R$ to be determined at $x$ by its pullbacks.
-Note that
-$i^{*}R=(i^{*}\nabla)\Phi|_{M}-i^{*}\sigma=(i^{*}\nabla)\phi-\sigma|_{M}=0$
-by the first equation in system (1).
-The transversality then yields that $R(\eta)$ vanishes at the intersection points of the fibers of $f$ with $M$.
-Since $R(\eta)$ is constant along the fibers and vanishes on $M$, it vanishes identically.
-Since $R(\xi)=R(\eta)=0$, we conclude that $R=0$, and hence the equation $\nabla\Phi=\sigma+R=\sigma$ holds on $f^{-1}(V)$.
-Because $\Phi$ satisfies $\nabla\Phi=\sigma$ globally on $f^{-1}(V)$, its local boundedness near the complement $X\setminus f^{-1}(V)$ follows from the holomorphic Poincaré lemma.
-Specifically, since $\nabla\sigma=0$, there exists a holomorphic local section s{s} satisfying $\nabla{s}=\sigma$ in a sufficiently small neighborhood of any point $x\in X\setminus f^{-1}(V)$.
-Consequently, we arrive at $\nabla(\Phi-{s})=0$ in the vicinity of $x$, which implies that $\Phi-{s}$ is a constant vector.
-The local boundedness of $\Phi$ is thus immediate.
-Consequently, the section $\Phi$ extends holomorphically across the proper analytic subset $X\setminus f^{-1}(V)$ to the entirety of $X$ by Riemann’s extension theorem, preserving the identity $\nabla\Phi=\sigma$ globally.
-∎
-Remark 2.2.
-The assumption of an $L^{2}$ fiber-wise solution makes this lemma particularly adaptable to geometric analysis settings, allowing one to rigidly extend weak distributional limits directly without establishing global uniform continuity beforehand.
+Since $fg=gh$, we deduce that $f(A)\subseteq A$. Thus, the endomorphisms $f,g,h$ restrict to the submodule $A$ and descend to the quotient $C=B/A$. We therefore obtain a short exact sequence
+$0\longrightarrow K(A)\longrightarrow K(B)\longrightarrow K(C)\longrightarrow 0$
+of complexes, which induces a long exact sequence
+$\dots\longrightarrow H^{i}(K(A))\longrightarrow H^{i}(K(B))\longrightarrow H^{i}(K(C))\longrightarrow H^{i+1}(K(A))\longrightarrow\cdots$ (1)
+of cohomology groups.
+We claim that the subcomplex $K(A)$ is acyclic.
+To verify exactness at the first term, we take $a\in A$ satisfying g(a)=0-g(a)=0.
+Since $g|_{A}$ is a monomorphism, we obtain $a=0$.
+Next, to establish exactness at the middle term, we take $(a,b)\in A\oplus A$ such that $f(a)+g(b)=0$.
+Because $g(A)=A$ and $g|_{A}$ is injective, in fact $g|_{A}$ is an automorphism.
+Consequently, there exists a unique $c\in A$ such that g(c)=a-g(c)=a.
+Applying $g$ to $h(c)$ yields $g(h(c))=(gh)(c)$, which becomes f(g(c))=f(a)=f(a)f(g(c))=f(-a)=-f(a) by virtue of the intertwining relation $fg=gh$.
+Since $f(a)+g(b)=0$, we have g(b)=f(a)=g(h(c))g(b)=-f(a)=g(h(c)).
+Because $g|_{A}$ is injective, we conclude $b=h(c)$.
+Thus, it holds that (a,b)=(g(c),h(c))(a,b)=(-g(c),h(c)), proving exactness at $A\oplus A$.
+Finally, to prove exactness at the terminal term, we choose an arbitrary $c\in A$.
+Since $g|_{A}$ is surjective, there exists $b\in A$ such that $g(b)=c$, demonstrating surjectivity.
+The acyclicity of $K(A)$ is therefore established.
+The exactness of sequence (1) then forces $H^{i}(K(B))=H^{i}(K(C))$ for all $i\in\mathbb{Z}$. ∎
+We now translate this abstract homological machinery into the geometric language of differential equations on compact Hermitian manifolds.
+By choosing our short exact sequence to reflect the Hodge decomposition of differential forms, we let the Dolbeault Laplacian assume the role of the intertwining operator.
+Because the Laplacian acts invertibly on the orthogonal complement of the harmonic forms, Lemma 2.1 establishes a dimension reduction: the defects of operators are completely captured by the finite-dimensional Dolbeault cohomology.
+Corollary 2.2.
+Let $X$ be a compact Hermitian manifold, and $L,A$ linear differential operators on $E:=\Omega^{p,q}(X)\cap\operatorname{Ker}(\bar{\partial})$.
+Suppose that $A\Delta_{\bar{\partial}}=\Delta_{\bar{\partial}}L$ and $L(\operatorname{Im}(\bar{\partial})\cap E)\subseteq\operatorname{Im}(\bar{\partial})$.
+Then, the following properties hold:
+  1. 1.
+    a solution $\omega\in E$ to $L\omega=\Delta_{\bar{\partial}}\omega=0$ is precisely a harmonic form of which Dolbeault cohomology class lies in the kernel of the induced operator $L^{*}\colon H^{p,q}(X)\rightarrow H^{p,q}(X)$;
+  2. 2.
+    for any $\alpha,\beta\in E$ satisfying $A\alpha=\Delta_{\bar{\partial}}\beta$, there exists $\omega\in E$ simultaneously solving $\Delta_{\bar{\partial}}\omega=\alpha$ and $L\omega=\beta$ if and only if the cohomology class $\alpha]=0$ and $\beta]\in\operatorname{Im}(L^{*})$.
+Proof.
+Consider the short exact sequence of vector spaces
+$0\longrightarrow\operatorname{Im}(\bar{\partial})\cap E\longrightarrow E\longrightarrow H^{p,q}(X)\longrightarrow 0.$
+Since the Dolbeault Laplacian $\Delta_{\bar{\partial}}$ commutes with $\bar{\partial}$, it restricts to an endomorphism on $E$.
+By Hodge theory, we have $\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$.
+Furthermore, as Green’s operator commutes with $\bar{\partial}$, the restriction of $\Delta_{\bar{\partial}}$ to the subspace $\operatorname{Im}(\bar{\partial})\cap E$ is an automorphism.
+Moreover, for any $\alpha\in\operatorname{Im}(\bar{\partial})\cap E$, there exists $\omega\in\operatorname{Im}(\bar{\partial})\cap E$ such that $\alpha=\Delta_{\bar{\partial}}\omega$.
+Therefore, we obtain that
+$A\alpha=A\Delta_{\bar{\partial}}\omega=\Delta_{\bar{\partial}}(L\omega).$
+Since $L\omega\in E$, we have $\Delta_{\bar{\partial}}(L\omega)\in\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$, which guarantees that the operator $A$ also preserves the subspace $\operatorname{Im}(\bar{\partial})\cap E$.
+The hypotheses $A\Delta_{\bar{\partial}}=\Delta_{\bar{\partial}}L$ and $L(\operatorname{Im}(\bar{\partial})\cap E)\subseteq\operatorname{Im}(\bar{\partial})$ thus ensure that the operators under consideration satisfy all premises of Lemma 2.1.
+Using the notations in the proof of Lemma 2.1, we have the isomorphism $H^{i}(K(E))=H^{i}(K(H^{p,q}(X)))$ for all $i\in\mathbb{Z}$.
+Because $\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$, the induced operator of $\Delta_{\bar{\partial}}$ on the quotient $H^{p,q}(X)$ is identically zero.
+The complex $K(H^{p,q}(X))$ thus simplifies to
+[…]
+Now, straightforward computation yields $H^{0}(K(H^{p,q}(X)))=\operatorname{Ker}(L^{*})$ and
+$H^{1}(K(H^{p,q}(X)))=\operatorname{Ker}(A^{*})\oplus\operatorname{Coker}(L^{*}),$
+which immediately translate to the two desired properties. ∎
 3 A Lemma on Kobayashi Hyperbolic Manifolds
@@
 Recall from birational geometry that the exceptional locus of a proper birational morphism between smooth projective varieties is uniruled, which once again contradicts the hyperbolicity of $X$.
-Quod erat demonstrandum. ∎
+∎
 Acknowledgement
@@
 Statements and Declarations
-The author certifies that the author has no affiliations with or involvement in any other organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
-This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author and the source, provide a link to the Creative Commons licence, and indicate if changes were made.
-The images or other third party material in this article are included in the article’s Creative Commons licence, unless indicated otherwise in a credit line to the material.
-If material is not included in the article’s Creative Commons licence and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder.
 Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.

### 22 Apr 2026 v3

@@
-A Note on Two Useful Lemmas in Complex Algebraic Geometry
+On Holomorphicity of Fiberwise Analytic Mappings
 Abstract
-We establish two useful lemmas in complex geometry.
-First, we investigate the homological properties of differential operators intertwined by the Dolbeault Laplacian on compact Hermitian manifolds.
-By utilizing the mapping cone of Koszul complexes, we prove that the analytical obstructions to solving specific overdetermined PDE systems reduce entirely to the finite-dimensional Dolbeault cohomology.
+We establish two rigidity results for mappings in complex geometry.
+First, we investigate the elevation of continuous mappings to holomorphic ones in the setting of proper fibrations.
+By leveraging relative morphism spaces and the Noetherian property of their local rings, we prove that a continuous, fiber-wise holomorphic map uniquely extends to a global holomorphic map, provided it is anchored by a sequence of holomorphic sections with mutually disjoint images.
+This result is then applied to establish the rigidity of homomorphisms between elliptic fibrations and Abelian schemes.
 Second, for a compact Kobayashi hyperbolic manifold fibered over $\mathbb{P}^{1}$, we prove that a continuous, fiber-wise holomorphic map of degree $1$ to a projective variety is a bi-holomorphic isomorphism, provided it is injective on a very ample hypersurface.
-Keywords: Dolbeault cohomology, Koszul complex, Complex manifold, Hyperbolicity
+Keywords: Complex manifold, Holomorphic mapping, Relative morphism space, Elliptic fibration, Kobayashi hyperbolicity, Birational geometry
 1 Introduction and Background
@@
 is holomorphic. Then $f$ itself is holomorphic.
-In accordance with the spirit of rigidity exemplified by Hartogs’ theorem, this paper investigates how rigid global structures in complex geometry can be enforced by partial algebraic or homological constraints.
-We divide our exploration into two distinct but philosophically related results.
-In Section 2, we turn our attention to differential operators on compact Hermitian manifolds.
-In the study of integrable systems, operators intertwined by a Laplacian exhibit highly constrained solution spaces.
-We abstract this phenomenon using homological algebra.
-By establishing an acyclicity lemma for intertwining chain complexes, we demonstrate how the Dolbeault Laplacian acts as an invertible intertwining operator on the space of exact forms.
-This allows us to prove that the infinite-dimensional analytical obstructions to solving certain coupled PDE systems collapse entirely into the finite-dimensional Dolbeault cohomology.
-In Section 3, we transition from the rigidity of differential equations to the rigidity of mappings.
-While Hartogs’ theorem relies on boundary integrals within bounded domains, transitioning to compact projective varieties requires global algebraic geometry.
-We present a rigidity lemma demonstrating how continuous, fiber-wise analytic mappings can be automatically elevated to global bi-holomorphic isomorphisms when anchored by a transverse algebraic intersection, specifically a very ample hypersurface, on a Kobayashi hyperbolic manifold.
-2 A Lemma on Holomorphic Differential Equations
-We begin our investigation by formulating a purely algebraic statement.
-By localizing the homological complexity to a specific submodule where an intertwining operator acts as an isomorphism, we equate the cohomology of the total space with that of the quotient.
+In accordance with the spirit of rigidity exemplified by Hartogs’ theorem, this paper investigates how global structures in complex geometry can be enforced by partial algebraic constraints.
+We divide our exploration into two distinct but related results concerning the rigidity of mappings.
+In both settings, we demonstrate that the presence of a positive algebraic object, specifically an infinite sequence of disjoint sections or a very ample divisor, acts as a geometric anchor.
+When combined with appropriate local analytic conditions, this anchor elevates partial data to global holomorphic maps or global bi-holomorphic isomorphisms.
+In Section 2, we turn our attention to the global extension of fiber-wise holomorphic mappings.
+While Hartogs’ theorem guarantees the extension of holomorphic functions across boundaries, elevating a continuous, fiber-wise holomorphic mapping to a globally holomorphic one on a proper fibration requires navigating the global geometry of the ambient space.
+We demonstrate that by anchoring the mapping to a sequence of holomorphic sections with mutually disjoint images, one can exploit the Noetherian nature of the local rings of the relative morphism space.
+This algebraic stabilization forces the mapping to be globally holomorphic, a rigidity result we apply to homomorphisms between elliptic fibrations and Abelian schemes.
+In Section 3, we transition to the rigidity of proper holomorphic mappings.
+We present a rigidity theorem demonstrating how continuous, fiber-wise analytic mappings can be automatically elevated to global bi-holomorphic isomorphisms when anchored by a transverse algebraic intersection, specifically a very ample hypersurface, on a Kobayashi hyperbolic manifold.
+2 Propagation of Holomorphicity on Proper Fibrations
+We begin by investigating the conditions under which a continuous mapping, known to be holomorphic on the individual fibers of a proper fibration, can be elevated to a global holomorphic mapping.
+The strategy relies on translating the pointwise agreement of sections into an infinite intersection of complex analytic subsets within the relative morphism space.
+By invoking the Noetherian property of analytic local rings, this descending chain must stabilize, enforcing global holomorphicity.
 Lemma 2.1.
-Let $0\rightarrow A\rightarrow B\rightarrow C\rightarrow 0$ be a short exact sequence of $R$-modules, and let $f,g,h\in\operatorname{End}_{R}(B)$ satisfy $fg=gh$, so that the sequence $K(B)$ of $R$-modules given by
-[…]
-is a chain complex.
-Suppose that $h(A)\subseteq g(A)=A$ and $g|_{A}$ is a monomorphism.
-Then $0\rightarrow A\rightarrow A^{\oplus 2}\rightarrow A\rightarrow 0$ forms an acyclic subcomplex $K(A)$ of $K(B)$, such that the cohomology groups of $K(C):=K(B)/K(A)$ are isomorphic to that of $K(B)$.
+Let $X,Y,Z$ be complex manifolds with $\dim(Y)=\dim(X)+1$ and $\phi\colon Y\rightarrow X$ a surjective proper holomorphic map with connected fibers.
+Let $\{s_{i}\}_{i\geq 1}$ be a sequence of holomorphic sections of $\phi$ with mutually disjoint images, and let $f\colon Y\rightarrow Z$ be a continuous mapping.
+Assume that $f|_{\phi^{-1}(x)}$ and $f\circ s_{i}$ are holomorphic for every $x\in X$ and all $i\geq 1$.
+Then $f$ is holomorphic.
 Proof.
-Since $fg=gh$, we deduce that $f(A)\subseteq A$. Thus, the endomorphisms $f,g,h$ restrict to the submodule $A$ and descend to the quotient $C=B/A$. We therefore obtain a short exact sequence
-$0\longrightarrow K(A)\longrightarrow K(B)\longrightarrow K(C)\longrightarrow 0$
-of complexes, which induces a long exact sequence
-$\dots\longrightarrow H^{i}(K(A))\longrightarrow H^{i}(K(B))\longrightarrow H^{i}(K(C))\longrightarrow H^{i+1}(K(A))\longrightarrow\cdots$ (1)
-of cohomology groups.
-We claim that the subcomplex $K(A)$ is acyclic.
-To verify exactness at the first term, we take $a\in A$ satisfying g(a)=0-g(a)=0.
-Since $g|_{A}$ is a monomorphism, we obtain $a=0$.
-Next, to establish exactness at the middle term, we take $(a,b)\in A\oplus A$ such that $f(a)+g(b)=0$.
-Because $g(A)=A$ and $g|_{A}$ is injective, in fact $g|_{A}$ is an automorphism.
-Consequently, there exists a unique $c\in A$ such that g(c)=a-g(c)=a.
-Applying $g$ to $h(c)$ yields $g(h(c))=(gh)(c)$, which becomes f(g(c))=f(a)=f(a)f(g(c))=f(-a)=-f(a) by virtue of the intertwining relation $fg=gh$.
-Since $f(a)+g(b)=0$, we have g(b)=f(a)=g(h(c))g(b)=-f(a)=g(h(c)).
-Because $g|_{A}$ is injective, we conclude $b=h(c)$.
-Thus, it holds that (a,b)=(g(c),h(c))(a,b)=(-g(c),h(c)), proving exactness at $A\oplus A$.
-Finally, to prove exactness at the terminal term, we choose an arbitrary $c\in A$.
-Since $g|_{A}$ is surjective, there exists $b\in A$ such that $g(b)=c$, demonstrating surjectivity.
-The acyclicity of $K(A)$ is therefore established.
-The exactness of sequence (1) then forces $H^{i}(K(B))=H^{i}(K(C))$ for all $i\in\mathbb{Z}$. ∎
-We now translate this abstract homological machinery into the geometric language of differential equations on compact Hermitian manifolds.
-By choosing our short exact sequence to reflect the Hodge decomposition of differential forms, we let the Dolbeault Laplacian assume the role of the intertwining operator.
-Because the Laplacian acts invertibly on the orthogonal complement of the harmonic forms, Lemma 2.1 establishes a dimension reduction: the defects of operators are completely captured by the finite-dimensional Dolbeault cohomology.
+Let $U$ be a Zariski dense open subset of $X$ consisting of regular values of $\phi$.
+For any $x\in U$, the fiber $Y_{x}:=\phi^{-1}(x)$ is a compact Riemann surface.
+Since $Y_{x}$ is compact, the infinite set of points $\{s_{i}(x)\}_{i\geq 1}$ possesses an accumulation point in $Y_{x}$.
+By the analytic continuation principle for complex manifolds, any two holomorphic mappings from $Y_{x}$ to $Z$ that coincide on $\{s_{i}(x)\}_{i\geq 1}$ are identical.
+Consider the relative morphism space
+$\pi\colon\operatorname{Hom}_{U}(\phi^{-1}(U),Z\times U):=\{(x,g):g\in\operatorname{Hom}(Y_{x},Z),x\in U\}\rightarrow U,\quad(x,g)\mapsto x$
+over $U$, of which connected components are complex analytic spaces by Douady’s theorem [2].
+Let $\sigma\colon U\rightarrow\operatorname{Hom}_{U}(\phi^{-1}(U),Z\times U)$ be the continuous section defined by $\sigma(x):=(x,f_{x})$, and denote by $\mathcal{H}$ the connected component of $\operatorname{Hom}_{U}(\phi^{-1}(U),Z\times U)$ containing $\sigma(U)$.
+For each integer $k\geq 1$, we define a holomorphic mapping $\Phi_{k}\colon\mathcal{H}\rightarrow(Z\times\cdots\times Z)\times U$ by
+$\Phi_{k}(x,g):=(g(s_{1}(x)),\dots,g(s_{k}(x)),x),$
+and define
+$A_{k}:=\Phi_{k}^{-1}(\{(f(s_{1}(x)),\dots,f(s_{k}(x)),x):x\in U\}).$
+Then $\{A_{k}\}_{k\geq 1}$ is a descending chain of complex analytic subsets in $\mathcal{H}$. By the analytic continuation principle, it holds that
+$\bigcap_{k=1}^{\infty}A_{k}=\sigma(U).$
+Because the local rings of $\mathcal{H}$ are Noetherian, the descending chain $\{A_{k}\}_{k\geq 1}$ must stabilize in a neighborhood of any point in $\sigma(U)$.
+Consequently, the restriction $\pi|_{\sigma(U)}$ is a bijective holomorphic mapping onto its image.
+Since $U$ is smooth, Zariski’s main theorem [4] immediately yields that $\sigma$ is holomorphic.
+By the universal property of the relative morphism space, the mapping $f|_{\phi^{-1}(U)}$ itself is holomorphic.
+Since $f$ is continuous and the singular locus $Y-\phi^{-1}(U)$ is a proper analytic subset of $Y$, Riemann’s extension theorem applied locally in the coordinate charts of $Z$ guarantees that $f$ is holomorphic on the entirety of $Y$.
+∎
+The application of Lemma 2.1 converts algebraic data, namely the stabilization of descending analytic chains, into global analytic rigidity.
+We now apply this machinery to elevate continuous homomorphisms between elliptic fibrations and Abelian schemes into holomorphic morphisms.
+By utilizing the group law of the elliptic fibration, a single torsion-free holomorphic section generates the required sequence of disjoint sections.
 Corollary 2.2.
-Let $X$ be a compact Hermitian manifold, and $L,A$ linear differential operators on $E:=\Omega^{p,q}(X)\cap\operatorname{Ker}(\bar{\partial})$.
-Suppose that $A\Delta_{\bar{\partial}}=\Delta_{\bar{\partial}}L$ and $L(\operatorname{Im}(\bar{\partial})\cap E)\subseteq\operatorname{Im}(\bar{\partial})$.
-Then, the following properties hold:
-  1. 1.
-    a solution $\omega\in E$ to $L\omega=\Delta_{\bar{\partial}}\omega=0$ is precisely a harmonic form of which Dolbeault cohomology class lies in the kernel of the induced operator $L^{*}\colon H^{p,q}(X)\rightarrow H^{p,q}(X)$;
-  2. 2.
-    for any $\alpha,\beta\in E$ satisfying $A\alpha=\Delta_{\bar{\partial}}\beta$, there exists $\omega\in E$ simultaneously solving $\Delta_{\bar{\partial}}\omega=\alpha$ and $L\omega=\beta$ if and only if the cohomology class $\alpha]=0$ and $\beta]\in\operatorname{Im}(L^{*})$.
+Let $\pi\colon E\rightarrow B$ be a smooth elliptic fibration and $\varphi\colon A\rightarrow B$ an Abelian scheme over a complex manifold $B$.
+Let $f\colon E\rightarrow A$ be a continuous mapping such that $\pi=\varphi\circ f$ and
+$f_{b}\colon\pi^{-1}(b)\rightarrow\varphi^{-1}(b),\quad P\mapsto f(P)$
+is a homomorphism for every $b\in B$.
+Assume there exists a holomorphic section $\sigma$ of $\pi$ such that $f\circ\sigma$ is holomorphic and $\sigma(b)$ is torsion-free for every $b\in B$.
+Then $f$ is holomorphic.
 Proof.
-Consider the short exact sequence of vector spaces
-$0\longrightarrow\operatorname{Im}(\bar{\partial})\cap E\longrightarrow E\longrightarrow H^{p,q}(X)\longrightarrow 0.$
-Since the Dolbeault Laplacian $\Delta_{\bar{\partial}}$ commutes with $\bar{\partial}$, it restricts to an endomorphism on $E$.
-By Hodge theory, we have $\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$.
-Furthermore, as Green’s operator commutes with $\bar{\partial}$, the restriction of $\Delta_{\bar{\partial}}$ to the subspace $\operatorname{Im}(\bar{\partial})\cap E$ is an automorphism.
-Moreover, for any $\alpha\in\operatorname{Im}(\bar{\partial})\cap E$, there exists $\omega\in\operatorname{Im}(\bar{\partial})\cap E$ such that $\alpha=\Delta_{\bar{\partial}}\omega$.
-Therefore, we obtain that
-$A\alpha=A\Delta_{\bar{\partial}}\omega=\Delta_{\bar{\partial}}(L\omega).$
-Since $L\omega\in E$, we have $\Delta_{\bar{\partial}}(L\omega)\in\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$, which guarantees that the operator $A$ also preserves the subspace $\operatorname{Im}(\bar{\partial})\cap E$.
-The hypotheses $A\Delta_{\bar{\partial}}=\Delta_{\bar{\partial}}L$ and $L(\operatorname{Im}(\bar{\partial})\cap E)\subseteq\operatorname{Im}(\bar{\partial})$ thus ensure that the operators under consideration satisfy all premises of Lemma 2.1.
-Using the notations in the proof of Lemma 2.1, we have the isomorphism $H^{i}(K(E))=H^{i}(K(H^{p,q}(X)))$ for all $i\in\mathbb{Z}$.
-Because $\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$, the induced operator of $\Delta_{\bar{\partial}}$ on the quotient $H^{p,q}(X)$ is identically zero.
-The complex $K(H^{p,q}(X))$ thus simplifies to
-[…]
-Now, straightforward computation yields $H^{0}(K(H^{p,q}(X)))=\operatorname{Ker}(L^{*})$ and
-$H^{1}(K(H^{p,q}(X)))=\operatorname{Ker}(A^{*})\oplus\operatorname{Coker}(L^{*}),$
-which immediately translate to the two desired properties. ∎
-3 A Lemma on Kobayashi Hyperbolic Manifolds
+For each integer $k\geq 1$, we define a section $s_{k}\colon B\rightarrow E$ by $s_{k}(b):=k\sigma(b)$, where the multiplication is taken with respect to the group structure of the elliptic fibration $\pi$.
+Since $\sigma(b)$ is torsion-free for every $b\in B$, the sequence of sections $\{s_{k}\}_{k\geq 1}$ has mutually disjoint images.
+Since $f$ is continuous, we have that $f_{b}$ is also continuous for all $b\in B$.
+But any continuous homomorphism between complex Lie groups is holomorphic, ensuring that $f|_{\pi^{-1}(b)}$ is holomorphic for all $b\in B$.
+Because the restriction $f_{b}\colon\pi^{-1}(b)\rightarrow\varphi^{-1}(b)$ is a group homomorphism for every $b\in B$, we obtain
+$(f\circ s_{k})(b)=f_{b}(k\sigma(b))=kf_{b}(\sigma(b))=k(f\circ\sigma)(b)$
+for all $b\in B$.
+Therefore we have the identity $f\circ s_{k}=\mu_{k}\circ(f\circ\sigma)$, where $\mu_{k}\colon A\rightarrow A$ denotes the integer multiplication morphism on the smooth Abelian scheme $A$.
+Since $A$ is Abelian, the mapping $\mu_{k}$ is holomorphic.
+By the assumption that $f\circ\sigma$ is holomorphic, the composition $f\circ s_{k}$ is holomorphic for every integer $k\geq 1$.
+The desired result then follows immediately from Lemma 2.1. ∎
+Remark 2.3.
+The proof of Corollary 2.2 encapsulates the overarching theme of this section: rigid global structures in complex geometry can be rigorously enforced by combining proper fibrations with the positive constraint of a torsion-free section.
+The algebraic stabilization within the relative morphism space serves as the essential bridge between continuous topological data and complex analytic rigidity.
+3 Rigidity of Mappings from Kobayashi Hyperbolic Manifolds
 We now transition from functions to mappings.
@@
 We divide the proof into three main steps: Establishing that the restrictions of $\phi$ on the fibers are finite, then proving their images are mutually disjoint, and finally upgrading the continuous map $\phi$ to a biholomorphism.
-Step 1: Since $H\in\operatorname{div}(X)$ is ample, the complex manifold $X$ is projective by Kodaira’s embedding theorem. Let
+Step 1: Since $H\in\operatorname{div}(X)$ is ample, the complex manifold $X$ is projective by Kodaira’s embedding theorem [3]. Let
 $\Sigma:=\mathbb{P}^{1}-\{b_{1},\dots,b_{m}\}$
@@
 For any $t\in\Sigma$, the fiber $X^{t}:=f^{-1}(t)$ is a smooth projective subvariety of $X$ of pure dimension n1n-1.
-Since $\phi^{t}:=\phi|_{X^{t}}$ is holomorphic and $Y$ is projective, Remmert’s proper mapping theorem and Chow’s theorem ensure that the image $Y^{t}:=\phi(X^{t})$ is a projective subvariety of $Y$.
+Since $\phi^{t}:=\phi|_{X^{t}}$ is holomorphic and $Y$ is projective, Remmert’s proper mapping [4] theorem and Chow’s theorem [3] ensure that the image $Y^{t}:=\phi(X^{t})$ is a projective subvariety of $Y$.
 We first establish the injectivity of the restriction $\phi^{t}$ for any fixed $t\in\Sigma$.
@@
 Because $H$ is a very ample divisor, intersecting a projective curve $C$ at a single point with a multiplicity of exactly 1 forces $C$ to be rational, that is, bi-holomorphic to $\mathbb{P}^{1}$.
-However, as $X$ is Kobayashi hyperbolic, there exists no non-constant holomorphic mapping from $\mathbb{P}^{1}$ into $X$ by Brody’s theorem on hyperbolicity.
+However, as $X$ is Kobayashi hyperbolic, there exists no non-constant holomorphic mapping from $\mathbb{P}^{1}$ into $X$ by Brody’s theorem [1] on hyperbolicity.
 This is a contradiction.
@@
 Now, recall from birational geometry that the exceptional locus of a proper birational morphism from a smooth projective variety is uniruled, meaning it is covered by rational curves.
-However, as $X$ is Kobayashi hyperbolic, by Brody’s theorem, it contains no rational curves, which leads to a contradiction.
+However, as $X$ is Kobayashi hyperbolic, by Brody’s theorem [1], it contains no rational curves, which leads to a contradiction.
 This proves that $\phi^{t}$ is indeed a finite morphism.
@@
 for all $t\in\mathbb{P}^{1}$.
-Since $Y$ is a smooth projective variety and $g\colon Y\rightarrow\mathbb{P}^{1}$ is a holomorphic fibration, Bertini’s theorem ensures that the general fiber of $g$ is smooth, and therefore normal.
-Thus, for a generic $t\in\Sigma$, the finite morphism $\phi^{t}\colon X^{t}\rightarrow Y^{t}$ of degree 1 onto a normal variety is an isomorphism by Zariski’s main theorem.
+Since $Y$ is a smooth projective variety and $g\colon Y\rightarrow\mathbb{P}^{1}$ is a holomorphic fibration, Bertini’s theorem [6] ensures that the general fiber of $g$ is smooth, and therefore normal.
+Thus, for a generic $t\in\Sigma$, the finite morphism $\phi^{t}\colon X^{t}\rightarrow Y^{t}$ of degree 1 onto a normal variety is an isomorphism by Zariski’s main theorem [6].
 This in particular proves that $\phi$ is injective almost everywhere.
@@
 Moreover, any continuous section of an unbranched analytic covering is necessarily holomorphic, so $\phi^{*}$ is holomorphic on $\Sigma$.
-Because $\phi$ is holomorphic along fibers and the section $\phi^{*}$ is also holomorphic, Osgood’s lemma on separate holomorphicity guarantees that $\phi$ is holomorphic on $f^{-1}(\Sigma)$.
+Because $\phi$ is holomorphic along fibers and the section $\phi^{*}$ is also holomorphic, Osgood’s lemma [4] on separate holomorphicity guarantees that $\phi$ is holomorphic on $f^{-1}(\Sigma)$.
 Since $\phi$ is continuous and the singular locus $X-f^{-1}(\Sigma)$ forms a proper algebraic subset, Riemann’s extension theorem guarantees that $\phi$ is holomorphic on the entirety of $X$.
@@
 Recall from birational geometry that the exceptional locus of a proper birational morphism between smooth projective varieties is uniruled, which once again contradicts the hyperbolicity of $X$.
-∎
+Quod erat demonstrandum. ∎
+Remark 3.2.
+Theorem 3.1 illustrates another facet of mapping regularity in complex geometry.
+While the results in Section 2 rely on an infinite sequence of disjoint sections to propagate holomorphicity, this theorem demonstrates that a single very ample hypersurface provides sufficient algebraic anchoring to elevate a continuous mapping to a global bi-holomorphic isomorphism.
+The intrinsic absence of rational curves in the Kobayashi hyperbolic manifold prevents the formation of exceptional loci, ensuring that the induced birational morphism is necessarily an isomorphism.
 Acknowledgement
-The author is deeply grateful to Weiyi Zhang for very useful and inspiring discussions.
+The author is deeply grateful to Weiyi Zhang for very useful and inspiring discussions. The author expresses gratitude to the reviewers for various suggestions.
 This research was completed while the author was studying at the Mathematics Institute of the University of Warwick.
 The author therefore would like to thank the University of Warwick for its hospitality.
+The author is in debt to the Russian mathematical society, especially, to all the professors teaching the math in Moscow program, and most importantly, to the author’s supervisor Alexander Petrovich Veselov at Loughborough University, as they cultivated the author’s mathematical literacy and maturity.
 Funding Information
@@
 Statements and Declarations
+The author certifies that the author has no affiliations with or involvement in any other organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
+This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author and the source, provide a link to the Creative Commons licence, and indicate if changes were made.
+The images or other third party material in this article are included in the article’s Creative Commons licence, unless indicated otherwise in a credit line to the material.
+If material is not included in the article’s Creative Commons licence and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder.
 Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.